OpenStudy (anonymous):

At a certain college the ratio to men to women is 6 to 5. If there are 1500 men, how many women are there

OpenStudy (apoorvk):

let the total no. of women be x. then, (no. of men)/(no.of women) = 6/5 so, 1500/x = 6/5 solve and enjoy!!

OpenStudy (anonymous):

I took 1500 divided by 5 and got 300 is that right?

OpenStudy (apoorvk):

nah..... why 1500 divided by 3?? see my solution.. you gotta equate 'ratio' with no. of men to women!

OpenStudy (anonymous):

i dont understand what you did though

OpenStudy (anonymous):

So 1500 divided by 5 then take 6 divided by 5 thats is what i m getting at

OpenStudy (apoorvk):

see. the question says. ratio of men to women is 6 to 5. it means that for every 6 men, we find 5 women. so now, if there are 1500 men, we find how many women then?? we know ratio is constant, i.e. 6 to 5 so i made an equation. clearer?

OpenStudy (apoorvk):

can you solve the equation i made above in my first post??

OpenStudy (anonymous):

Still dont understand

OpenStudy (apoorvk):

okay.. i ll go slow. just answer me. okay.. suppose you teacher tells you that for every two question you answer correctly, she gives you 3 dollars. you find that you have 15 dollars in the end. how many do you think you must have had answered correctly??

OpenStudy (apoorvk):

confused eh?